# compat=1.1.0

### v1

@@ Appendix A Details of FOPT and Gravitational Waves
Here, $\Upsilon=1-\frac{1}{\sqrt{1+2\tau_{sw}\mathcal{H}_{*}}}$ is a suppression factor Guo:2020grp where $\tau_{\rm sw}\sim R_{*}/\bar{U}_{f}$, mean bubble separation, $R_{*}=(8\pi)^{1/3}v_{w}\beta$ and rms fluid velocity, $\bar{U}_{f}=\sqrt{3\kappa_{sw}\alpha_{*}/4(1+\alpha_{*})}$; $z_{p}\sim 10$. The spectrum generated by the turbulence in the plasma is given by Caprini:2015zlo ; Athron:2023xlk ; Caprini:2024hue
@@
Figure 6: Left panel: Variation of DM-nucleon elastic scattering cross section with DM mass for different choices of the kinetic mixing parameter. Right panel: Same as the left panel but for DM-electron scattering. For both the panels, $m_{Z_{D}}$=8×$\times 10^{-3}$ GeV and $g_{D}$=0.6.
@@ Appendix B Direct detection of light DM
The kinetic mixing of Z-${\rm Z}_{D}$ made possible to probe dark matter through elastic scattering with nucleons in direct detection experiments like CRESST III CRESST:2019jnq , XENON1T XENON:2019gfn etc. The spin independent DM-nucleon cross section for the experiment target nucleus with atomic number Z and mass number A can be written Borah:2021pet as
$\sigma_{\chi N}^{SI}=\frac{g^{2}g_{D}^{2}\epsilon^{2}\mu_{\chi N}^{2}}{\pi m_{Z_{D}}^{4}}\frac{(Zf_{p}+(A-Z)f_{n})^{2}}{A^{2}}$ (30)
where, $\mu_{\chi N}=\frac{m_{\chi}m_{N}}{(m_{\chi}+m_{N})}$ is the reduced mass of DM-nucleon, $\epsilon$ is Z-ZD kinetic mixing, $f_{p}$ is interaction strength of proton and $f_{n}$ is interaction strength of neutron.
However, the dark matter electron scattering in direct detection experiments XENON1T XENON:2019gfn , PandaX-4T PandaX:2022xqx can also provide bound on DM mass. The DM-electron reference cross section Essig:2015cda ; Prabhu:2022dtm is given by
$\overline{\sigma}_{e}=\frac{16\pi\alpha\alpha_{D}\epsilon^{2}\mu_{\chi e}^{2}}{m_{Z_{D}}^{4}}$ (31)
where, $\alpha$ is fine structure constant, $\alpha_{D}=g_{D}^{2}/4\pi$, $\mu_{\chi e}=\frac{m_{\chi}m_{e}}{(m_{\chi}+m_{e})}$ is reduced mass of DM-electron. The variation of DM-nucleon and DM-electron scattering cross section with DM mass in the range of our interest is shown in Fig. 6.

### v2

@@ Appendix A Details of FOPT and Gravitational Waves
Here, $\Upsilon=1-\frac{1}{\sqrt{1+2\tau_{sw}\mathcal{H}_{*}}}$ is a suppression factor Guo:2020grp where $\tau_{\rm sw}\sim R_{*}/\bar{U}_{f}$, mean bubble separation, $R_{*}=(8\pi)^{1/3}v_{w}\beta^{-1}$ and rms fluid velocity, $\bar{U}_{f}=\sqrt{3\kappa_{sw}\alpha_{*}/4(1+\alpha_{*})}$; $z_{p}\sim 10$. The spectrum generated by the turbulence in the plasma is given by Caprini:2015zlo ; Athron:2023xlk ; Caprini:2024hue
@@
Figure 9: Left panel: Variation of spin-independent DM-nucleon elastic scattering cross section with DM mass for different choices of the singlet-Higgs mixing parameter. Right panel: Same as the left panel but for DM-electron scattering. For both the panels, $m_{\Phi_{1}}=10$ MeV and $\lambda_{\chi\Phi_{1}}$=0.01.
@@ Appendix B Direct detection of light DM
DM $\chi$, can interact with nucleon due to Higgs portal interactions facilitated by the mixing between $\Phi_{1}$ and the SM Higgs $h$, parameterized by $\sin{\theta_{{\Phi_{1}}h}}$. The spin independent DM-nucleon cross section for a target nucleus with atomic number $Z$ and mass number $A$ can be written as Adhikary:2024btd
$\sigma_{\chi N}^{SI}=\frac{\mu^{2}_{{\chi}N}}{4\pi A^{2}}\left(Zf_{p}+(A-Z)f_{n}\right)^{2},$ (46)
where $\mu_{{\chi}N}=\frac{m_{\chi}m_{N}}{(m_{\chi}+m_{N})}$ is the reduced mass of DM-nucleon system and the interaction strengths $f_{p}$, $f_{n}$ of proton and neutron with DM are given as
$f_{p,n}=\sum_{q=u,d,s}f^{p,n}_{T_{q}}\alpha_{q}\frac{m_{p,n}}{m_{q}}+\sum_{q=c,t,b}f^{p,n}_{TG}\alpha_{q}\frac{m_{p,n}}{m_{q}},$ (47)
where $f^{p}_{T_{u}}=0.018$, $f^{p}_{T_{d}}=0.027$, $f^{p}_{T_{s}}=0.037$, $f^{n}_{T_{u}}=0.013$, $f^{n}_{T_{d}}=0.040$, $f^{n}_{T_{s}}=0.037$ and $f^{p}_{TG}=1-f^{p}_{T_{u}}-f^{p}_{T_{d}}-f^{p}_{T_{s}}$, $f^{n}_{TG}=1-f^{n}_{T_{u}}-f^{n}_{T_{d}}-f^{n}_{T_{s}}$ Ellis:2018dmb . The $\alpha_{q}$ is defined by
$\alpha_{q}=\lambda_{\chi\Phi_{1}}\sin{\theta}_{\Phi_{1}h}\frac{m_{q}}{v}\left(\frac{1}{m^{2}_{\Phi_{1}}}-\frac{1}{m^{2}_{h}}\right),$ (48)
with $m_{h}\approx 125$ GeV being the mass of SM Higgs.
Moreover, the dark matter electron scattering in direct detection experiments XENON1T XENON:2019gfn , PandaX-4T PandaX:2022xqx can also provide bound on DM mass. The DM-electron cross section is given by Adhikary:2024btd ; Essig:2015cda ; Prabhu:2022dtm
$\sigma_{e}\approx\frac{16\pi\mu_{{\chi}e}\alpha\alpha_{\chi\Phi_{1}}\sin^{2}{\theta_{{\Phi_{1}}h}}}{m^{4}_{\Phi_{1}}}|F(q)|^{2},$ (49)
with $\alpha_{\chi\Phi_{1}}=\lambda^{2}_{\chi\Phi_{1}}/4\pi$ and $\mu_{\chi e}=\frac{m_{\chi}m_{e}}{m_{\chi}+m_{e}}$ refers the reduced mass of DM-electron system. $\alpha$ denotes the fine structure constant, and $F(q)=1$ for a massive mediator. The variations of DM-nucleon and DM-electron scattering cross section with DM mass in the range of our interest are shown in Fig. 9.
Figure 10: Comparison of conversion rate of $\Phi_{2}\rightarrow\Phi^{\dagger}_{2}$ for different value of $\lambda_{12}$ and other parameters same as BP1.
Appendix C Efficient conversion of $\Phi_{2}\rightarrow\Phi^{\dagger}_{2}$
The decay of asymmetric $\chi$ can also create asymmetry in $\Phi_{2}$, but due to the scalar coupling $\lambda_{12}(\Phi_{2}\Phi_{2}\Phi_{1}\Phi_{1}+{\rm h.c.})$, efficient conversion of $\Phi_{2}\rightarrow\Phi^{\dagger}_{2}$ can happen before and during FOPT, effectively washing out the asymmetry stored $\Phi_{2}$. The cross-section for this process is $\sigma_{\rm conv}=\lambda_{12}^{2}/(16\pi s)$. The corresponding thermal averaged cross section Cannoni:2015wba can be written as
$\langle\sigma v\rangle_{\rm conv}=\frac{1}{8Tm_{\Phi_{1}}^{2}m_{\Phi_{2}}^{2}K_{2}(m_{\Phi_{1}}/T)K_{2}(m_{\Phi_{2}}/T)}\int_{(m_{\Phi_{1}}+m_{\Phi_{2}})^{2}}^{\infty}ds\frac{\lambda(s,m_{\Phi_{1}},m_{\Phi_{2}})}{\sqrt{s}}K_{1}(\frac{\sqrt{s}}{T})\sigma_{\rm conv}$ (50)
where, the Mandelstam triangular function $\lambda(s,m_{1},m_{2})=(s-(m_{1}+m_{2})^{2})(s-(m_{1}-m_{2})^{2})$. Fig. 10 shows the comparison of this conversion rate with Hubble for different value of $\lambda_{12}$, and other parameter same as benchmark point BP1 in table 1. Clearly, the process remains in equilibrium around the phase transition epoch, efficiently washing out the asymmetry stored in $\Phi_{2}$.